OpenStudy (electrochika):

dy/dx of y= (x + sqrt x)^-2

OpenStudy (anonymous):

chain rule for this one

OpenStudy (anonymous):

\[\frac{d}{dx}(x+\sqrt{x})^{-2}=-2(x+\sqrt{x})^{-3}\times (1+\frac{1}{2\sqrt{x}})\]

OpenStudy (electrochika):

I don't know how to get there

OpenStudy (anonymous):

ok it is the chain rule, the power rule, and the fact that i have memorized \[\frac{d}{dx}\sqrt{x}=\frac{1}{2\sqrt{x}}\]

OpenStudy (electrochika):

I know that you bring -2 out front, but where does the ^-3 come from? or the second half? I've missed a week of school.

OpenStudy (anonymous):

you have a composite function. i looks like \[(f(x))^{-2}\] were \[f(x)=x+\sqrt{x}\] the derivative of something to the power of minus 2 is minus 2 something to the power of -3 (by the power rule) times the derivative of "something"

OpenStudy (anonymous):

\[\frac{d}{dx}f(x)^n=nf^{n-1}(x)\times f'(x)\]

OpenStudy (anonymous):

just like \[\frac{d}{dx}x^n=nx^{n-1}\]

OpenStudy (anonymous):

so to answer more specifically, the -3 comes from -2 - 1

OpenStudy (electrochika):

ohh so the second half is just the derivative?

OpenStudy (anonymous):

the second half is the derivative of the "inside function" yes. and the derivative of \[x+\sqrt{x}\] is \[1+\frac{1}{2\sqrt{x}}\]

OpenStudy (electrochika):

okay so every time i have a \[\sqrt{x}\] , then it's \[1/2\sqrt{x}\]?